OpenStudy (anonymous):

In a multiple-choice test, each question has five options. Students will get 5 points for each correct answer; lose 2 point for each incorrect answer; and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show your calculations for expected value. Show your work for full credit.

OpenStudy (anonymous):

expected value. take \[5\times \frac{1}{5}-2\times \frac{4}{5}\]

OpenStudy (anonymous):

get \[-\frac{3}{5}\] so expected value is negative. if you guess on average you will lose \[-\frac{3}{5}=-.6\] of a point

OpenStudy (anonymous):

best to leave blank, don't lose anything

OpenStudy (anonymous):

Disadvantage.